# LLM-based NLG Evaluation: Current Status and Challenges

### v2

@@ 3.3 Task Instructions
In human evaluation, this part usually comes in the form of a task description or evaluation steps. They can also exist at the same time. The task description states the annotation in a more general way, and the evaluation steps, which can be considered as Chain-of-Thought, explicitly describe what to do at each step. In the context of prompting LLMs for NLG evaluation, the demonstrations used in few-shot prompting are also included.
Form and requirements. Several studies from an Eval4NLP 2023 shared task Leiter et al. (2023) have explored task instructions in various settings. Kim et al. (2023a) conduct experiments on different templates and lengths of task descriptions and evaluation steps. Kotonya et al. (2023) generate task instructions with LLMs or improve existing task instructions with LLMs. Additionally, He et al. (2023a) evaluate generative reasoning using LLMs by asking them first to generate their own answers, and then conduct a quantitative analysis of the text to be evaluated.
Analysis and explanations. LLMs can include analysis or explanation in their evaluations, which is a key point that distinguishes them from previous automatic evaluation metrics. Early explorations into prompting LLMs for NLG evaluation mostly do not examine the impact of whether LLMs are required to analyze and explain on evaluation results. However, Chiang and Lee (2023b) explore different types of evaluation instructions in summarization evaluation and dialogue evaluation, finding that explicitly asking large models to provide analysis or explanation achieve higher correlation with human judgments. Besides, the quality of the analysis and explanation generated by LLMs itself requires additional manual evaluation Leiter et al. (2023). Naismith et al. (2023) compare the explanations written by humans and generated by GPT-4 and conduct a simple corpus analysis on the generated explanations.
@@ 3.4 Input Content
The types of input content mainly depend on the evaluation criteria and are relatively fixed. For most task-specific evaluation criteria, such as the faithfulness of a summary Luo et al. (2023); Gao et al. (2023), the source document is needed in addition to the target text to be evaluated. For task-independent criteria, such as fluency Hu et al. (2023); Chiang and Lee (2023b), only the text to be evaluated needs to be provided, though many works also provide the source document Wang et al. (2023a); Liusie et al. (2023). Other types of input content can be provided as required by the specific task. Kocmi and Federmann (2023b) use two different settings when evaluating machine translation: providing references and not providing references. Guan et al. (2023) provide relevant facts and context when evaluating whether a text conforms to the facts. Exceptionally, Shu et al. (2023) add the output of other automatic evaluation metrics to the input of the LLM.
@@ 3.5 Evaluation Criteria
The evaluation targeting specific aspects is used in numerous studies of human evaluation for NLG, such as text summarization, story generation, dialogue, and text simplification. Evaluation criteria, i.e., the definitions of aspects are key in this context. Most evaluation criteria in LLM-based evaluation are directly derived from human evaluation. However, a few studies have attempted to let LLMs generate or improve evaluation criteria. Liu et al. (2023e) use a few human-rated examples as seeds to let LLMs draft some candidate evaluation criteria, and then further filter them based on the performance of LLMs using these criteria on a validation set, to obtain the final evaluation criteria. Kim et al. (2023c) designed an LLM-based interactive evaluation system, which involves using LLMs to review the evaluation criteria provided by users, including eliminating ambiguities in criteria, merging criteria with overlapping meanings, and decomposing overly broad criteria. Additionally, Ye et al. (2023a) propose a hierarchical aspect classification system with 12 subcategories, demonstrating that under the proposed fine-grained aspect definitions, human evaluation and LLM-based evaluation are highly correlated. Besides, the chain-of-aspects approach improves LLMs’ ability to evaluate on a specific aspect by having LLMs score on some related aspects before generating the final score Gong and Mao (2023).
@@ 3.6 Role and Interaction
We include in this section the evaluation strategies that either use the same LLMs in different ways or involve different LLMs. The former can be further divided into chain-style and network-style interactions.
Chain-style interaction. Inspired by human evaluators, Yuan et al. (2024) have LLMs score a batch of examples to be evaluated each time. Specifically, the evaluation process is divided into three stages: analysis, ranking, and scoring. Similar to QA-based evaluation metrics Durmus et al. (2020), Fu et al. (2023b) assess the faithfulness of summaries in two stages: treating LLMs as question generators to generate a question from the summary; then having LLMs answer the question using the source document. Differently, when Hu et al. (2023) use GPT-4 to evaluate the faithfulness of summaries, it first asks GPT-4 to extract event units from the summary, then verifies whether these event units meet the requirements, and finally judges whether the event units are faithful to the source document.
Network-style interaction. Unlike chain-style interactions, network-style interactions involve the dispersion and aggregation of information. In network-style interactions, LLMs on the same layer play similar roles. ChatEval Chan et al. (2023) is a framework for evaluating content through debates among multiple LLMs, with three communication strategies designed among the three types of LLMs: One-By-One, Simultaneous-Talk, and Simultaneous-Talk-with-Summarizer. Zhang et al. (2023b) find that under certain conditions, widening and deepening the network of LLMs can better align its evaluation with human judgments. Saha et al. (2023) propose a branch-solve-merge strategy, assigning LLMs the roles of decomposing problems, solving them, and aggregating answers, thereby improving the accuracy and reliability of evaluations. Wu et al. (2023) assume that different people such as politicians and the general public have different concerns about the quality of news summaries, use LLMs to play different roles in evaluation accordingly, and aggregate the results finally.
Different LLMs. Different from having the same LLM play different roles, some research has used different LLMs (such as GPT-4 and Claude) in their studies. In pairwise comparisons, previous work mostly used a single LLM as the evaluator, which may not be fair. In light of this, Bai et al. (2023) design a decentralized Peer-examination method, using different LLMs as evaluators and then aggregating the results. Further, Li et al. (2023c) let different LLMs serve as evaluators in pairwise comparisons and then have them go through a round of discussion to reach the final result. Additionally, Cohen et al. (2023) evaluate the factuality of texts through the interaction of two LLMs, where the LLM that generated the text acts as the examinee and the other LLM as the examiner.
@@ 3.7 Pros and Cons
The benefits of prompting LLMs for NLG evaluation are exciting. First, for the first time, people can express evaluation criteria and evaluation methods in natural language within the prompts given to LLMs, providing great flexibility. Where previously people needed to design specific evaluation metrics for different NLG tasks or even different aspects of a single task, now they only need to modify the prompts for LLMs. Secondly, surprisingly, LLMs have the ability to generate explanations while assessing texts, making this approach somewhat interpretable. Furthermore, in many NLG task, prompting LLMs for evaluation has achieved state-of-the-art correlations with human judgments.
However, as many studies have pointed out, this type of approach still has many limitations. Wang et al. (2023b) note that when using ChatGPT and GPT-4 for pairwise comparisons, the order of the two texts can affect the evaluation results, which is known as position bias. To alleviate this issue, Li et al. (2023d) propose a strategy of splitting, aligning, and then merging the two texts to be evaluated into the prompt. Also, LLM evaluators tend to favor longer, more verbose responses Zheng et al. (2023) and responses generated by themselves Liu et al. (2023a). Wu and Aji (2023) show that compared to answers that are too short or grammatically incorrect, answers with factual errors are considered better by LLMs. Liu et al. (2023d) demonstrate through adversarial meta-evaluation that LLMs without references are not suitable for evaluating dialogue responses in closed-ended scenarios: they tend to score highly on responses that conflict with the facts in the dialogue history. Zhang et al. (2023a) also present the robustness issues of LLMs in dialogue evaluation through adversarial perturbations. Shen et al. (2023) indicate that LLM evaluators have a lower correlation with human assessments when scoring high-quality summaries. In addition, Hada et al. (2023) state that LLM-based evaluators have a bias towards high scores, especially in non-Latin languages like Chinese and Japanese. Beyond these shortcomings of performance, both ChatGPT and GPT-4 are proprietary models, and their opacity could lead to irreproducible evaluation results.
@@ 4 Fine-tuning LLMs
4.1 Overview
As mentioned above, despite the exciting performance of prompting LLMs like ChatGPT and GPT-4 for NLG evaluation, several shortcomings in practice are inevitable, such as high costs, possibly irreproducible results, and potential biases in LLMs. In response, recent research has shifted towards fine-tuning smaller, open-source LLMs specifically for evaluation purposes, aiming to achieve performance close to GPT-4 in NLG evaluation. Representative works of this type include PandaLM Wang et al. (2023e), Prometheus Kim et al. (2023b), Shepherd Wang et al. (2023c), TIGERScore Jiang et al. (2023), INSTRUCTSCORE Xu et al. (2023), Auto-J Li et al. (2023a), CritiqueLLM Ke et al. (2023) and JudgeLM Zhu et al. (2023). Their main ideas are similar, involving the elaborate construction of high-quality evaluation data, followed by accordingly fine-tuning open-source base LLMs. Nevertheless, there are certain discrepancies in the designs across different works, such as the usage of references and evaluation criteria. We have summarized the key different components of these methods in Table 1 for comparison, which will be elaborated on next.
Method Data Construction Evaluation Method Base LLM Reference Required
Instruction Source Annotator Scale Result Mode Details Specific Criteria
PandaLM Alpaca 52K GPT-3.5 300K Comparison Reason & Reference Unified LLaMA 7B No
Prometheus GPT-4 Construction GPT-4 100K Scoring Reason Explicit LLaMA-2-Chat 7B & 13B Yes
Shepherd Community Critique Data & 9 NLP Tasks Data Human 1317 Overall Judgement Error Identifying & Refinement Unified LLaMA 7B No
TIGERScore 23 Distinctive Text Generation Datasets GPT-4 48K MQM Error Analysis Implicit LLaMA-2 7B & 13B No
INSTRUCTSCORE GPT-4 Construction GPT-4 40K MQM Error Analysis Implicit LLaMA 7B Yes
AUTO-J Real-world User Queries from Preference Datasets GPT-4 4396 Scoring & Comparison Reason Implicit LLaMA-2-Chat 13B No
CritiqueLLM AlignBench & ChatGPT Augmentation GPT-4 9332 Scoring Reason Unified ChatGLM-2 6B, 12B & 66B Flexible
JudgeLM GPT4All-LAION, ShareGPT Alpaca-GPT4 & Dolly-15K GPT-4 100K Scoring & Comparison Reason Unified Vicuna 7B, 13B & 33B Flexible
Table 1: Comparison of the different key components among the representative methods of fine-tuning LLMs.
@@ 4.2 Data Construction
The primary differences in these works actually lie in the construction of instructions, with the purpose of improving either diversity or reliability for the better generalization ability of the fine-tuned model. PandaLM and JudgeLM entirely sample from common instruction datasets, such as Alpaca 52K, while CritiqueLLM adopts small-scale sampling followed by ChatGPT augmentation. In contrast, Prometheus and INSTRUCTSCORE rely on GPT-4 to generate all the instructions based on seed data, whereas Auto-J and Shepherd use real-world data. Moreover, since large-scale human annotation is impractical, most works utilize GPT-4 as the powerful annotator, except for PandaLM and Shepherd, which use GPT-3.5 and human annotation on small-scale community data, respectively. During the construction, they basically all design detailed prompts or guidance and apply heuristic filtering strategies and post-processing methods to mitigate noise. Overall, despite the possible higher quality of human annotation, the corresponding drawback is the difficulty in constructing large-scale datasets, which in turn may hinder adequate model training, while using LLMs for construction is the opposite situation.
@@ 4.3 Evaluation Method
As with prompting LLMs, the evaluation methods adopted in these works are highly diversified, involving different evaluation criteria, result modes, and usages of the reference. Given that current instruction-response scenarios encompass different types of tasks, it is unsuitable to specify unified evaluation criteria as in traditional NLG tasks. However, some works still do it this way, while some other methods let LLM annotators adaptively and implicitly reflect the required criteria in their evaluations, like PandaLM, TIGERScore, and AUTO-J. In particular, AUTO-J has meticulously crafted 332 evaluation criteria, matched to different tasks. Furthermore, Prometheus explicitly incorporates evaluation criteria into the inputs of the model, expecting flexible evaluation based on various customized criteria.

### v3

@@ 3.2 Task Instructions
In human evaluation, task instruction usually comes in the form of a task description or evaluation steps. They can also exist at the same time. The task description states the annotation in a more general way, and the evaluation steps, which can be considered as Chain-of-Thought, explicitly describe what to do at each step. In the context of prompting LLMs for NLG evaluation, we discuss three broad categories of influences: various templates of prompts (Leiter et al., 2023; Kim et al., 2023; Kotonya et al., 2023; He, Zhang, and Roth, 2024), in-context examples (Jain et al., 2023; Kotonya et al., 2023; Hasanbeig et al., 2023), and whether LLMs are required to provide analyses or explanations (Chiang and Lee, 2023b; Naismith, Mulcaire, and Burstein, 2023).
Form and requirements. Several studies from an Eval4NLP 2023 shared task (Leiter et al., 2023) have explored task instructions in various settings. For example, Kim et al. (2023) conduct experiments on different templates and lengths of task descriptions and evaluation steps, finding that providing clear and straightforward instructions, akin to those explained to humans, is more effective. Kotonya et al. (2023) generate task instructions with LLMs or improve existing task instructions with LLMs. Moreover, Leiter and Eger (2024) conduct a larger-scale prompt exploration for the evaluation of machine translation and summarization based on the Eval4NLP 2023 shared task. Somewhat differently, He, Zhang, and Roth (2024) evaluate generative reasoning using LLMs by asking them first to generate their own answers, and then conduct a quantitative analysis of the text to be evaluated. Additionally, explicit evaluation requirements and output formats are typically included in the instructions, and the evaluation results are extracted using regular expression matching. Early LLMs may sometimes provide unrecognizable evaluation results or refuse to conduct evaluation due to their limited instruction-following capabilities (Gao et al., 2023). This issue can be mitigated through multiple sampling or setting random outputs, and it basically does not exist in the currently more advanced and powerful LLMs.
Analysis and explanations. LLMs are able to include analysis or explanation in their evaluations, which is a key point that distinguishes them from previous automatic evaluation metrics. Early explorations into prompting LLMs for NLG evaluation mostly do not examine the impact of whether LLMs are required to analyze and explain on evaluation results. However, Chiang and Lee (2023b) explore different types of evaluation instructions in summarization evaluation and dialogue evaluation, finding that explicitly asking large models to provide analysis or explanation achieve higher correlation with human judgments. Besides, the quality of the analysis and explanation generated by LLMs itself requires additional manual evaluation (Leiter et al., 2023). Naismith, Mulcaire, and Burstein (2023) compare the explanations written by humans and generated by GPT-4 and conduct a simple corpus analysis on the generated explanations, finding that GPT-4 has strong potential to produce ratings that are comparable to human ratings on discourse coherence, accompanied by clear rationales.
@@ 3.3 Input Content
The types of input content mainly depend on the evaluation criteria and are relatively fixed. For most task-specific evaluation criteria, such as the faithfulness of a summary (Luo, Xie, and Ananiadou, 2023; Gao et al., 2023), the source document is needed in addition to the target text to be evaluated. For task-independent criteria, such as fluency (Hu et al., 2023; Chiang and Lee, 2023b), only the text to be evaluated needs to be provided, though many works also provide the source document (Wang et al., 2023a; Liusie, Manakul, and Gales, 2024). Other types of input content can be provided as required by the specific task. Kocmi and Federmann (2023b) use two different settings when evaluating machine translation: providing references and not providing references and find that GPT-4 without references can also outperform all existing reference-based metrics. Guan et al. (2024) provide relevant facts and context when evaluating whether a text conforms to the facts. Exceptionally, Shu et al. (2024) add the output of other automatic evaluation metrics to the input of the LLM.
@@ 3.4 Evaluation Criteria
The evaluation targeting specific aspects is used in numerous studies of human evaluation for NLG, such as text summarization, story generation, dialogue, and text simplification. Evaluation criteria, i.e., the definitions of aspects are key in this context. Most evaluation criteria in LLM-based evaluation are directly derived from human evaluation. However, a few studies have attempted to let LLMs generate or improve evaluation criteria. Liu et al. (2024c) use a few human-rated examples as seeds to let LLMs draft some candidate evaluation criteria, and then further filter them based on the performance of LLMs using these criteria on a validation set, to obtain the final evaluation criteria. Kim et al. (2024d) designed an LLM-based interactive evaluation system, which involves using LLMs to review the evaluation criteria provided by users, including eliminating ambiguities in criteria, merging criteria with overlapping meanings, and decomposing overly broad criteria. Additionally, Ye et al. (2024) propose a hierarchical aspect classification system with 12 subcategories, demonstrating that under the proposed fine-grained aspect definitions, human evaluation and LLM-based evaluation are highly correlated. Besides, the chain-of-aspects approach improves LLMs’ ability to evaluate on a specific aspect by having LLMs score on some related aspects before generating the final score (Gong and Mao, 2023).
@@ 3.5 Role and Interaction
We include in this section the evaluation strategies that either use the same LLMs in different ways or involve different LLMs (Bai et al., 2023; Li, Patel, and Du, 2023; Cohen et al., 2023). The former can be further divided into chain-style (Yuan et al., 2024; Fu et al., 2023; Hu et al., 2023) and network-style interactions (Chan et al., 2024; Zhang et al., 2023; Saha et al., 2024; Wu et al., 2023).
Chain-style interaction. Inspired by human evaluators, Yuan et al. (2024) have LLMs score a batch of examples to be evaluated each time. Specifically, the evaluation process is divided into three stages: analysis, ranking, and scoring. Similar to QA-based evaluation metrics (Durmus, He, and Diab, 2020), Fu et al. (2023) assess the faithfulness of summaries in two stages: treating LLMs as question generators to generate a question from the summary; then having LLMs answer the question using the source document. Differently, when Hu et al. (2023) use GPT-4 to evaluate the faithfulness of summaries, it first asks GPT-4 to extract event units from the summary, then verifies whether these event units meet the requirements, and finally judges whether the event units are faithful to the source document.
Network-style interaction. Unlike chain-style interactions, network-style interactions involve the dispersion and aggregation of information. In network-style interactions, LLMs on the same layer play similar roles. ChatEval (Chan et al., 2024) is a framework for evaluating content through debates among multiple LLMs, with three communication strategies designed among the three types of LLMs: One-By-One, Simultaneous-Talk, and Simultaneous-Talk-with-Summarizer. Zhang et al. (2023) find that under certain conditions, widening and deepening the network of LLMs can better align its evaluation with human judgments. Saha et al. (2024) propose a branch-solve-merge strategy, assigning LLMs the roles of decomposing problems, solving them, and aggregating answers, thereby improving the accuracy and reliability of evaluations. Wu et al. (2023) assume that different people such as politicians and the general public have different concerns about the quality of news summaries, use LLMs to play different roles in evaluation accordingly, and aggregate the results finally.
Different LLMs. Different from having the same LLM play different roles, some research has used different LLMs (such as GPT-4 and Claude) in their studies. The use of a single LLM as evaluator may introduce bias, resulting in unfair evaluation results. In light of this, Bai et al. (2023) design a decentralized Peer-examination method, using different LLMs as evaluators and then aggregating the results. Further, Li, Patel, and Du (2023) let different LLMs serve as evaluators in pairwise comparisons and then have them go through a round of discussion to reach the final result. Additionally, Cohen et al. (2023) evaluate the factuality of texts through the interaction of two LLMs, where the LLM that generated the text acts as the examinee and the other LLM as the examiner.
@@ 3.6 Pros and Cons
The benefits of prompting LLMs for NLG evaluation are exciting. First, for the first time, people can express evaluation criteria and evaluation methods in natural language within the prompts given to LLMs, providing great flexibility. Where previously people needed to design specific evaluation metrics for different NLG tasks or even different aspects of a single task, now they only need to modify the prompts for LLMs. Secondly, surprisingly, LLMs have the ability to generate explanations while assessing texts, making this approach somewhat interpretable. Furthermore, in many NLG tasks, prompting LLMs for evaluation has achieved state-of-the-art correlations with human judgments.
However, as many studies have pointed out, this type of approach still has many limitations. Wang et al. (2024a) note that when using ChatGPT and GPT-4 for pairwise comparisons, the order of the two texts can affect the evaluation results, which is known as position bias. To alleviate this issue, Li et al. (2023b) propose a strategy of splitting, aligning, and then merging the two texts to be evaluated into the prompt. Also, LLM evaluators tend to favor longer, more verbose responses (Zheng et al., 2023) and responses generated by themselves (Liu et al., 2023a). Wu and Aji (2025) show that compared to answers that are too short or grammatically incorrect, answers with factual errors are considered better by LLMs. Liu et al. (2023b) demonstrate through adversarial meta-evaluation that LLMs without references are not suitable for evaluating dialogue responses in closed-ended scenarios: they tend to score highly on responses that conflict with the facts in the dialogue history. Zhang et al. (2024) also present the robustness issues of LLMs in dialogue evaluation through adversarial perturbations. Shen et al. (2023) indicate that LLM evaluators have a lower correlation with human assessments when scoring high-quality summaries. In addition, Hada et al. (2024b) state that LLM-based evaluators have a bias towards high scores, especially in non-Latin languages like Chinese and Japanese. Bavaresco et al. (2024) find that the performance of LLM-based evaluators exhibits significant variance depending on the dataset, evaluation criteria, and whether the evaluated texts are human-generated. Beyond these shortcomings of performance, both ChatGPT and GPT-4 are proprietary models, and their opacity could lead to irreproducible evaluation results.
@@ 4 Fine-tuning LLMs
As mentioned above, despite the exciting performance of prompting LLMs like ChatGPT and GPT-4 for NLG evaluation, several shortcomings in practice are inevitable, such as high costs, possibly irreproducible results, and potential biases in LLMs. In response, recent research has shifted towards fine-tuning smaller, open-source LLMs specifically for evaluation purposes, aiming to achieve performance close to GPT-4 in NLG evaluation. Representative works of this type include PandaLM (Wang et al., 2024c), Prometheus (Kim et al., 2024a), Prometheus 2 (Kim et al., 2024c), Shepherd (Wang et al., 2023b), TIGERScore (Jiang et al., 2024), INSTRUCTSCORE (Xu et al., 2023), Auto-J (Li et al., 2024), CritiqueLLM (Ke et al., 2023), JudgeLM (Zhu, Wang, and Wang, 2023), Themis (Hu et al., 2024), CompassJudger-1 (Cao et al., 2024) and Self-Taught (Wang et al., 2024b). Their main ideas are similar, involving the elaborate construction of high-quality evaluation data, followed by fine-tuning open-source foundation LLMs with specific methods. Nevertheless, there are certain discrepancies in the designs across different works, such as the usage of references and evaluation criteria. We have summarized the key different components of these methods in Table 2 and Table 3 for comparison, which will be elaborated in the following sections.
Method Data Construction Foundation LLM
Instruction Source Annotator Scale
PandaLM Alpaca 52K GPT-3.5 300K LLaMA 7B
Prometheus GPT-4 Construction GPT-4 100K LLaMA-2-Chat 7B & 13B
Prometheus 2 FEEDBACK COLLECTION GPT-4 200K Mistral-7B Mixtral-8x7B
Shepherd Community Critique Data & 9 NLP Tasks Data Human 1317 LLaMA 7B
TIGERScore 23 Distinctive Text Generation Datasets GPT-4 48K LLaMA-2 7B & 13B
INSTRUCTSCORE GPT-4 Construction GPT-4 40K LLaMA 7B
AUTO-J Real-world User Queries from Preference Datasets GPT-4 4396 LLaMA-2-Chat 13B
CritiqueLLM AlignBench & ChatGPT Augmentation GPT-4 9332 ChatGLM-2 6B, 12B & 66B
JudgeLM GPT4All-LAION, ShareGPT Alpaca-GPT4 & Dolly-15K GPT-4 100K Vicuna 7B, 13B & 33B
Themis NLG-Eval with 58 NLG Evaluation Datasets Human & GPT-4 67K Llama-3-8B
Self-Taught Screened WildChat Llama-3-70B 20K Llama-3-70B
CompassJudger-1 Sampling from existing datasets Mixture 900K Qwen-2.5 1.5B, 7B, 14B & 32B
Table 2: Comparison of the different key components among the representative methods of fine-tuning LLMs (Part 1).
Method Evaluation Method Reference Required
Result Mode Details Specific Criteria
PandaLM Comparison Reason & Reference Unified No
Prometheus Scoring Reason Explicit Yes
Prometheus 2 Scoring & Comparison Reason Explicit Yes
Shepherd Overall Judgement Error Identifying & Refinement Unified No
TIGERScore MQM Error Analysis Implicit No
INSTRUCTSCORE MQM Error Analysis Implicit Yes
AUTO-J Scoring & Comparison Reason Implicit No
CritiqueLLM Scoring Reason Unified Flexible
JudgeLM Scoring & Comparison Reason Unified Flexible
Themis Scoring Reason Explicit No
Self-Taught Comparison Reason Unified No
CompassJudger-1 Scoring & Comparison Reason Explicit No
Table 3: Comparison of the different key components among the representative methods of fine-tuning LLMs (Part 2).
@@ 4.1 Data Construction
The primary differences in these works actually lie in the construction of instructions, with the purpose of improving either diversity or reliability for the better generalization ability of the fine-tuned model. PandaLM and JudgeLM entirely sample from common instruction datasets, such as Alpaca 52K, while CritiqueLLM adopts small-scale sampling followed by ChatGPT augmentation. In contrast, Prometheus and INSTRUCTSCORE rely on GPT-4 to generate all the instructions based on seed data, whereas Auto-J and Shepherd use real-world data. Moreover, since large-scale human annotation is impractical, most works utilize GPT-4 as the powerful annotator, except for PandaLM and Shepherd, which use GPT-3.5 and human annotation on small-scale community data, respectively. Specifically, Themis focuses on NLG tasks and combines existing human evaluations with additional evaluations from GPT-4, selecting more consistent training data. Self-Taught uses the evaluation results from the model to fine-tune itself (Llama-3-70B), considering it already possesses strong capabilities. During the construction, these studies basically all design detailed prompts or guidance and apply heuristic filtering strategies and post-processing methods to mitigate noise. Overall, despite the possible higher quality of human annotation, the corresponding drawback is the difficulty in constructing large-scale datasets, which in turn may hinder adequate model training, while using LLMs for construction is the opposite situation.
@@ 4.2 Evaluation Method
As with prompting LLMs, the evaluation methods adopted in these works are highly diversified, involving different evaluation criteria, result modes, and usages of the reference. Given that current instruction-response scenarios encompass different types of tasks, it is unsuitable to specify unified evaluation criteria as in traditional NLG tasks. However, some works still do it this way, while some other methods let LLM annotators adaptively and implicitly reflect the required criteria in their evaluations, like PandaLM, TIGERScore, and AUTO-J. In particular, AUTO-J has meticulously crafted 332 evaluation criteria, matched to different tasks. Furthermore, Prometheus and Themis explicitly incorporate evaluation criteria into the evaluation instructions, and CompassJudger can work either with or without evaluation criteria, enabling flexible evaluation based on various customized criteria.